OpenStudy (anonymous):

"a" pens cost "b" dollars, how many dollars would "c" pens cost at the same rate?

OpenStudy (texaschic101):

a/b = c/x --- (a pens to b dollars = c pens to x dollars) cross multiply (a)(x) = (b)(c) ax = bc x = (bc)/a c pens cost (bc)/a dollars